what is the force required for a steel wire to double its length when its area of cross section is 1 sqaure cm and young modulus is 200 Gpa, as the length of the wire is double, the change in length is equal to its original length


1
Expert's answer
2021-02-18T18:55:46-0500
"Y=\frac{Fl}{A\Delta l}\\Y=\frac{Fl}{Al}\\200\cdot10^9=\frac{F}{10^{-4}}\\F=2\cdot10^{7}\ N"